# Forecasting realized volatility in the stock market: a path-dependent perspective

### v1

@@ 4.3 In-sample fitting results
    Note: The adjusted $R^{2}$ value denotes the adjusted goodness-of-fit. The symbols *, **∗ ∗, and ***∗ ∗ ∗ denote the rejection of the null hypothesis at the $10\%$, $5\%$, and $1\%$ significance levels, respectively. The values in parentheses correspond to the standard errors of the respective parameters, with all estimated results rounded to three decimal places.
@@
    Note: The adjusted $R^{2}$ value denotes the adjusted goodness-of-fit. The symbols *, **∗ ∗, and ***∗ ∗ ∗ denote the rejection of the null hypothesis at the $10\%$, $5\%$, and $1\%$ significance levels, respectively. The values in parentheses correspond to the standard errors of the respective parameters, with all estimated results rounded to three decimal places.
To prevent overfitting and further explain the predictive performance of the HAR-PD model family with path dependence, the LASSO method, as described in Section 3.2.1, was employed to construct the HAR-PD model under LASSO [^3] The specific code can be found at https://github.com/wangxiaobo018/realized-volatility.
@@
$\overline{R}_{2,t-5}$ $-5.437^{***}$ $(1.382)$
$\mathrm{{PDREQ}}_{t-1}^{+}$ $-1.388^{***}$ $(0.228)$
@@
$\mathrm{\overline{{PDREQ}}}_{t-5}^{-}$ $-1.900^{***}$ $(0.194)$
$\mathrm{\overline{{PDREQ}}}_{t-22}^{-}$ $-1.388^{**}$ $(0.228)$
@@
    Note: The adjusted $R^{2}$ value denotes the adjusted coefficient of determination. The symbols *, **∗ ∗, and ***∗ ∗ ∗ denote the rejection of the null hypothesis at the $10\%$, $5\%$, and $1\%$ significance levels, respectively. The values in parentheses correspond to the standard errors of the respective parameters, with all estimated results rounded to three decimal places.
Table 3 presents the statistical description of volatility based on five-minute high-frequency data, with values of all variables except Ljung-Box(10) test rounded to three decimal places. The distribution of volatility exhibits significant right skewness, indicating a higher frequency of large volatility values. Additionally, the kurtosis value is 95.914, suggesting a highly peaked distribution with fat tails, which implies the potential presence of extreme values. The ADF test statistics for all variables are negative, and the null hypothesis of a unit root is rejected at the $1\%$ significance level, indicating that all time series variables are stationary. Taking the SSE as an example, Tables 4 and 5 present the parameter estimation results of the HAR and HAR-PD model families. In terms of adjusted $R^{2}$, the HAR-RS model demonstrates the highest adjusted goodness-of-fit among the HAR family models, which is consistent with the volatility prediction model findings reported by Zhang and Zhang (2023). As shown in Table 5, the HAR-PD-RV model exhibits the adjusted $R^{2}$ of 0.530, with parameters $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$ being statistically significant at the $1\%$ level, substantially outperforming the baseline HAR-RV model’s adjusted $R^{2}$ of 0.406. A notable improvement in the adjusted $R^{2}$ is also observed when comparing the HAR-CJ model with the HAR-PD-CJ model. The HAR-PD model family demonstrates higher significance levels compared to the baseline HAR model family. The HAR-PD-RS model achieves the highest adjusted $R^{2}$ with multiple significant parameters. Similarly, the empirical quantile decomposition-based HAR-PD-REQ model shows superior the adjusted $R^{2}$ compared to the baseline HAR-REX model. Overall, the HAR-PD model family consistently demonstrates higher adjusted $R^{2}$ values than the HAR model family. Table 6 presents the model variables selected through LASSO, where variables $R_{2,t-1}$ and $\overline{R}_{2,t-5}$ were consistently selected across all models in the LASSO-HAR-PD family, further substantiating the contribution of volatility features to the models. The information criterion of the HAR model family and the LASSO-HAR-PD model family shown in the Table 7 shows that a balance has been achieved between the complexity and fit of the LASSO-HAR-PD model family. The improved model does not achieve a better fit by increasing complexity. And the HAR-REQ model constructed by empirical quantiles has smaller AIC and BIC than the HAR-REX model.
@@
HAR-RV $-64062.49$ $-64030.30$
HAR-CJ $-61909.70$ $-61858.94$
HAR-RS $-63022.55$ $-62971.09$
HAR-REX $-62253.74$ $-62182.94$
HAR-REQ $-62468.68$ $-62193.92$
HAR-PD-RV $-62403.08$ $-62370.92$
LASSO-HAR-PD-CJ $-62986.16$ $-62947.54$
LASSO-HAR-PD-RS $-63449.58$ $-63410.97$
LASSO-HAR-PD-REX $-63008.93$ $-62951.00$
LASSO-HAR-PD-REQ $-62823.72$ $-62759.39$
@@ 4.4.1 Rolling time window prediction and Out-of-sample prediction evaluation methods
    The sample data are divided into estimation sample (trian sample) and forecast sample (test sample). Following prior studies (Haugom et al., 2014; Feng et al., 2024), forecast sample the length of the forecast sample is generally 300, 600, 1000. In this paper, we will use the forecast sample of length 1000 in the robustness test. After data cleaning, the estimation sample is a fixed window of 4016 days from February 2, 2005, to July 6, 2021, and the forecast sample is the last 600 trading days.
@@
    Data from $t=1,2,\cdots,N=4295$ was selected as the total sample. In this case, when predicting 1-step into the future, with a fixed window of 4,016 observations ($H=4016$), the predicted sample is the last 600 days ($M=600$), similarly when predicting 5-step into the future, the predicted sample is 596 days ($M=596$), and when predicting 22-step into the future, the predicted sample is 579 days ($M=579$).
@@
To comprehensively validate the advantages of the path-dependent structure beyond merely comparing the HAR model family, several researchers (Moreno-Pino and Zohren, 2024; Hansen and Lunde, 2005) have incorporated additional prediction models as baseline comparisons. Thus, in this context, we utilize the GARCH model (Bollerslev, 1986) and its extensions as benchmark models.
Table 8: Out-of-sample forecast loss value of SSE for the future 1-step
@@
GARCH $1.629\times 10^{-8}$ $7.912\times 10^{-5}$ $10.468$ $1.863$ $-8.364$
EGARCH $1.076\times 10^{-8}$ $6.422\times 10^{-5}$ $4.666$ $1.411$ $-8.545$
FGARCH $8.008\times 10^{-9}$ $4.967\times 10^{-5}$ $2.135$ $0.972$ $-8.578$
TGARCH $8.652\times 10^{-9}$ $5.605\times 10^{-5}$ $2.511$ $1.116$ $-7.741$
HAR-RV $1.284\times 10^{-8}$ $5.773\times 10^{-5}$ $3.658$ $0.951$ $-8.414$
HAR-CJ $1.373\times 10^{-8}$ $6.976\times 10^{-5}$ $8.679$ $1.631$ $-8.409$
HAR-RS $4.427\times 10^{-9}$ $3.339\times 10^{-5}$ $0.836$ $0.578$ $-8.724$
HAR-REX $5.081\times 10^{-9}$ $3.432\times 10^{-5}$ $0.587$ $0.552$ $-8.697$
HAR-REQ $5.056\times 10^{-9}$ $3.344\times 10^{-5}$ $0.526$ $0.524$ $-8.698$
HAR-PD-RV $6.858\times 10^{-9}$ $3.955\times 10^{-5}$ $0.802$ $0.574$ $-8.607$
HAR-PD-CJ $4.693\times 10^{-9}$ $3.324\times 10^{-5}$ $0.780$ $0.536$ $-8.726$
HAR-PD-RS $\underline{\mathbf{4.185\times 10^{-9}}}$ $3.446\times 10^{-5}$ $1.035$ $0.625$ $-8.652$
HAR-PD-REX $4.923\times 10^{-9}$ $3.286\times 10^{-5}$ $0.522$ $0.511$ $-8.700$
HAR-PD-REQ $4.855\times 10^{-9}$ $\underline{\mathbf{3.116\times 10^{-5}}}$ $\underline{\mathbf{0.450}}$ $\underline{\mathbf{0.467}}$ $-8.707$
LASSO-HAR-PD-CJ $5.038\times 10^{-9}$ $3.277\times 10^{-5}$ $0.462$ $0.483$ $-8.701$
@@
LASSO-HAR-PD-REX $5.022\times 10^{-9}$ $3.463\times 10^{-5}$ $0.625$ $0.571$ $-8.692$
LASSO-HAR-PD-REQ $4.981\times 10^{-9}$ $3.317\times 10^{-5}$ $0.535$ $0.520$ $-8.698$
@@
When utilizing the aforementioned loss functions to evaluate model accuracy, relying solely on a single loss function for comparison is insufficient. If the prediction loss of one model is lower than that of another, it is directly concluded that the first model has higher predictive accuracy. However, this judgment is limited and cannot be easily generalized to other datasets or scenarios involving different loss functions. In particular, extreme outliers in the data may significantly impact the calculation of loss functions, potentially leading to misleading evaluations of the volatility model’s performance.
Building on this foundation, some scholars (White, 2000; Diebold and Mariano, 2002; Hansen and Lunde, 2005) have proposed different methods for evaluating the out-of-sample predictive ability of models. However, the most widely applied approach is the MCS method [^4] The specific MCS inspection process is shown in the A., introduced by Hansen et al. (2011)
@@ 4.4.2 Out-of-sample volatility forecasting
Owing to space limitations, the specific results see B. The MCS test is employed to select the optimal forecasting model at a given confidence level. The MCS test statistics, $T_{R}$ and $T_{max}$, along with the corresponding $p$-values, are obtained through 5000 bootstrap simulations. The MCS procedure is commonly utilized for evaluating sample models. Table 8, presents the optimal models under different loss functions after undergoing the MCS test. As shown in Table 8 the HAR-PD model family exhibits significantly improved predictive performance compared to the HAR model family when forecasting predicting 1-step. Based on the MCS test results, the following conclusions can be drawn:
@@
    As shown in the Table 8, we used CSI 300 as the dataset and used a rolling time window to predict the last 300 days of the dataset. And the predicted loss values of different models (including MSE, MAE, HMSE, HMAE, and QLIKE loss functions) are compared predicting 1-step into the future. In terms of loss values, the HAR-PD family of models performs well in all five loss functions, maintaining low loss values even under LASSO constraints. Notably, the HAR-PD-RS and LASSO-HAR-PD-RS models have significantly lower loss values than the other models for predicting the 1-step, 5-step, and 22-step predictions. In addition, the HAR-REQ model, which defines thresholds based on quantized values, also outperforms the HAR-REX model. The HAR-PD-REQ model has the lowest loss values under the MAE, HMSE and HMAE loss functions.
@@
    In comparing the HAR model family and the GARCH model family, found that the prediction accuracy of the HAR model family is higher than that of the GARCH model family, which is consistent with the findings of Luo and Chen (2020). Among the HAR model families, the HAR-RS model has the lowest loss value, while the HAR-CJ model has a better loss value than the HAR-RV model, which is consistent with the findings of Wen et al. (2016).
@@ 4.5 Robustness testing
In Section 4.4.1, a rolling window prediction was conducted using the last 600 observations as the prediction sample, following the methodology outlined in Yang et al. (2015). For robustness analysis, a shorter rolling window was selected compared to the out-of-sample forecast, which allowed for a longer prediction interval. To evaluate the model’s performance on actual unknown data, this study employed the first 3,616 observations as the rolling window length and the subsequent 1,000 observations as the prediction window. After data cleaning, the estimation sample spanned from January 18, 2005, to November 4, 2019, while the forecast sample covered the period from November 5, 2019, to December 24, 2023.
Secondly in order to assess the model’s forecasting ability on other data, new data will be used here for forecasting. Section 4.4.1 uses the SSE, while the robustness test uses 970 days of CSI 300 Index data from January 2, 2020 to December 29, 2023. The data is divided into a 670-day estimation sample from January 2, 2020 to October 11, 2022 and a 300-day forecast sample from October 12, 2022 to December 29, 2023.
Finally, to comprehensively evaluate the superiority of the extended model, additional baseline models were introduced for comparative analysis beyond the basic HAR model family, following the approaches of Qu et al. (2018) and Moreno-Pino and Zohren (2024), who incorporated various traditional baseline models to demonstrate the performance advantages of their extended models.

### v2

@@ 4.3 In-sample fitting results
    Note: The adjusted $R^{2}$ value denotes the adjusted goodness-of-fit. The symbols *, **, and *** denote the rejection of the null hypothesis at the $10\%$, $5\%$, and $1\%$ significance levels, respectively. The values in parentheses correspond to the standard errors of the respective parameters, with all estimated results rounded to three decimal places.
@@
    Note: The adjusted $R^{2}$ value denotes the adjusted goodness-of-fit. The symbols *, **, and *** denote the rejection of the null hypothesis at the $10\%$, $5\%$, and $1\%$ significance levels, respectively. The values in parentheses correspond to the standard errors of the respective parameters, with all estimated results rounded to three decimal places.
To prevent overfitting and further explain the predictive performance of the HAR-PD model family with path dependence, the LASSO method, as described in Section 3.2.1, was employed to construct the HAR-PD model under LASSO 44endnote: 4The specific code can be found at https://github.com/wangxiaobo018/realized-volatility.
@@
$\overline{R}_{2,t-5}$ $5.437^{***}$ $(1.382)$
$\mathrm{{PDREQ}}_{t-1}^{+}$ $1.388^{***}$ $(0.228)$
@@
$\mathrm{\overline{{PDREQ}}}_{t-5}^{-}$ $1.900^{***}$ $(0.194)$
$\mathrm{\overline{{PDREQ}}}_{t-22}^{-}$ $1.388^{**}$ $(0.228)$
@@
    Note: The adjusted $R^{2}$ value denotes the adjusted coefficient of determination. The symbols *, **, and *** denote the rejection of the null hypothesis at the $10\%$, $5\%$, and $1\%$ significance levels, respectively. The values in parentheses correspond to the standard errors of the respective parameters, with all estimated results rounded to three decimal places.
Table 3 presents the statistical description of volatility based on five-minute high-frequency data, with values of all variables except Ljung-Box(10) test rounded to three decimal places. The distribution of volatility exhibits significant right skewness, indicating a higher frequency of large volatility values. Additionally, the kurtosis value is 95.914, suggesting a highly peaked distribution with fat tails, which implies the potential presence of extreme values. The ADF test statistics for all variables are negative, and the null hypothesis of a unit root is rejected at the $1\%$ significance level, indicating that all time series variables are stationary. Taking the SSE as an example, Tables 4 and 5 present the parameter estimation results of the HAR and HAR-PD model families. In terms of adjusted $R^{2}$, the HAR-RS model demonstrates the highest adjusted goodness-of-fit among the HAR family models, which is consistent with the volatility prediction model findings reported by Zhang and Zhang (2023). As shown in Table 5, the HAR-PD-RV model exhibits the adjusted $R^{2}$ of 0.503, with parameters $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$ being statistically significant at the $1\%$ level, substantially outperforming the benchmark HAR-RV model’s adjusted $R^{2}$ of 0.406. A notable improvement in the adjusted $R^{2}$ is also observed when comparing the HAR-CJ model with the HAR-PD-CJ model. The HAR-PD model family demonstrates higher significance levels compared to the benchmark HAR model family. The HAR-PD-RS model achieves the highest adjusted $R^{2}$ with multiple significant parameters. Similarly, the empirical quantile decomposition-based HAR-PD-REQ model shows superior the adjusted $R^{2}$ compared to the benchmark HAR-REX model. Overall, the HAR-PD model family consistently demonstrates higher adjusted $R^{2}$ values than the HAR model family. Table 6 presents the model variables selected through LASSO, where variables $R_{2,t-1}$ and $\overline{R}_{2,t-5}$ were consistently selected across all models in the LASSO-HAR-PD family, further substantiating the contribution of volatility features to the models. The information criterion of the HAR model family and the LASSO-HAR-PD model family shown in the Table 7 shows that a balance has been achieved between the complexity and fit of the LASSO-HAR-PD model family. The improved model does not achieve a better fit by increasing complexity. And the HAR-REQ model constructed by empirical quantiles has smaller AIC and BIC than the HAR-REX model.
@@
HAR-RV 64062.49-64062.49 64030.30-64030.30
HAR-CJ 61909.70-61909.70 61858.94-61858.94
HAR-RS 63022.55-63022.55 62971.09-62971.09
HAR-REX 62253.74-62253.74 62182.94-62182.94
HAR-REQ 62468.68-62468.68 62193.92-62193.92
HAR-PD-RV 62403.08-62403.08 62370.92-62370.92
LASSO-HAR-PD-CJ 62986.16-62986.16 62947.54-62947.54
LASSO-HAR-PD-RS 63449.58-63449.58 63410.97-63410.97
LASSO-HAR-PD-REX 63008.93-63008.93 62951.00-62951.00
LASSO-HAR-PD-REQ 62823.72-62823.72 62759.39-62759.39
@@ 4.4.1 Rolling time window prediction and Out-of-sample prediction evaluation methods
    The sample data are divided into in-sample (trian sample) and out-of-sample (test sample). Following prior studies (Haugom et al., 2014; Feng et al., 2024), out-of-sample the length of the out-of-sample is generally 300, 600, 1000. In this paper, we will use the out-of-sample of length 1000 in the robustness test. After data cleaning, the in-sample is a fixed window of 4016 days from February 2, 2005, to July 6, 2021, and the out-of-sample is the last 600 trading days.
@@
    Data from $t=1,2,\cdots,N=4295$ was selected as the total sample. In this case, when predicting one-step-ahead into the future, with a fixed window of 4,016 observations ($H=4016$), the predicted sample is the last 600 days ($M=600$), similarly when predicting five-step-ahead into the future, the predicted sample is 596 days ($M=596$), and when predicting twenty-two-step-ahead into the future, the predicted sample is 579 days ($M=579$).
@@
To comprehensively validate the advantages of the path-dependent structure beyond merely comparing the HAR model family, several researchers (Moreno-Pino and Zohren, 2024; Hansen and Lunde, 2005) have incorporated additional prediction models as benchmark comparisons. Thus, in this context, we utilize the GARCH model (Bollerslev, 1986) and its extensions as benchmark models.
Table 8: Out-of-sample forecast loss value of SSE for the one-step-ahead predictions
@@
GARCH $6.442\times 10^{-9}$ $4.657\times 10^{-5}$ $1.975$ $0.937$ 8.608-8.608
EGARCH $5.692\times 10^{-9}$ $4.657\times 10^{-5}$ $1.975$ $0.937$ 8.629-8.629
TGARCH $5.259\times 10^{-9}$ $4.841\times 10^{-5}$ $1.942$ $0.995$ 8.657-8.657
GJRGARCH $8.954\times 10^{-9}$ $6.429\times 10^{-5}$ $4.726$ $1.553$ 8.602-8.602
HAR-RV $7.025\times 10^{-9}$ $4.887\times 10^{-5}$ $1.924$ $1.001$ 8.609-8.609
HAR-CJ $6.180\times 10^{-9}$ $4.133\times 10^{-5}$ $1.006$ $0.743$ 8.648-8.648
HAR-RS $4.427\times 10^{-9}$ $3.339\times 10^{-5}$ $0.836$ $0.578$ 8.724-8.724
HAR-REX $5.081\times 10^{-9}$ $3.432\times 10^{-5}$ $0.587$ $0.552$ 8.697-8.697
HAR-REQ $5.056\times 10^{-9}$ $3.344\times 10^{-5}$ $0.526$ $0.524$ 8.698-8.698
HAR-PD-RV $6.858\times 10^{-9}$ $3.955\times 10^{-5}$ $0.802$ $0.574$ 8.607-8.607
HAR-PD-CJ $4.693\times 10^{-9}$ $3.324\times 10^{-5}$ $0.780$ $0.536$ 8.726-8.726
HAR-PD-RS $\underline{\mathbf{4.185\times 10^{-9}}}$ $3.446\times 10^{-5}$ $1.035$ $0.625$ 8.652-8.652
HAR-PD-REX $4.923\times 10^{-9}$ $3.286\times 10^{-5}$ $0.522$ $0.511$ 8.700-8.700
HAR-PD-REQ $4.855\times 10^{-9}$ $\underline{\mathbf{3.116\times 10^{-5}}}$ $\underline{\mathbf{0.450}}$ $\underline{\mathbf{0.467}}$ 8.707-8.707
LASSO-HAR-PD-CJ $5.038\times 10^{-9}$ $3.277\times 10^{-5}$ $0.462$ $0.483$ 8.701-8.701
@@
LASSO-HAR-PD-REX $5.022\times 10^{-9}$ $3.463\times 10^{-5}$ $0.625$ $0.571$ 8.692-8.692
LASSO-HAR-PD-REQ $4.981\times 10^{-9}$ $3.317\times 10^{-5}$ $0.535$ $0.520$ 8.698-8.698
@@
When utilizing the aforementioned loss functions to evaluate model accuracy, relying solely on a single loss function for comparison is insufficient. If the prediction loss of one model is lower than that of another, it is directly concluded that the first model has higher predictive accuracy. However, this judgment is limited and cannot be easily generalized to other datasets or scenarios involving different loss functions. In particular, extreme outliers in the data may significantly impact the calculation of loss functions, potentially leading to misleading evaluations of the volatility model’s performance. Building on this foundation, some scholars (White, 2000; Diebold and Mariano, 2002; Hansen and Lunde, 2005) have proposed different methods for evaluating the out-of-sample predictive ability of models. However, the most widely applied approach is the MCS method 55endnote: 5The specific MCS inspection process is shown in the A., introduced by Hansen et al. (2011). The MCS test is employed to select the optimal forecasting model at a given confidence level. The MCS test statistics, $T_{R}$ and $T_{max}$, along with the corresponding $p$-values, are obtained through 5000 bootstrap simulations. In subsequent chapters, we will use the MCS test to test the accuracy of out-of-sample forecast results.
@@ 4.4.2 Out-of-sample volatility forecasting
Table 8 shows that with SSE as the dataset and a rolling time window, we predict the last 600 days of the dataset and present the one-step-ahead prediction loss values under various loss functions (MSE, MAE, HMSE, HMAE, and QLIKE). The HAR model family has much lower prediction loss values than the GARCH model family, which is consistent with Luo and Chen (2020). The HAR-PD model family performs well in all five loss functions and maintains low loss values even under the LASSO constraint. Notably, in Tables 11 and 12, when predicting the future five and twenty-two-step-ahead, the HAR-PD-RS and LASSO-HAR-PD-RS models have significantly lower loss values than the other models. In addition, the HAR-REQ model, which defines the threshold based on the empirical quantile, outperforms the HAR-REX model, and the HAR-PD-REQ model has the lowest loss values under the MAE, HMSE, and HMAE loss functions.
In order to further test the prediction accuracy of the model, according to the suggestion of Amendola et al. (2020); Hansen et al. (2003), we use the MCS test based on the QLIKE loss function and under the MSE loss function to evaluate the out-of-sample prediction accuracy. Based on the MCS test results 66endnote: 6Owing to space limitations, the specific results see B., the following conclusions can be drawn:
@@
    As shown in Tables 13, 14, and 15 for the MCS test results under the QLIKE and MSE loss functions, the HAR model family has more models passing the test than the GARCH model family. This is consistent with the findings of previous studies Vortelinos (2017); Qu et al. (2018); Gong and Lin (2019). Moreover, the HAR-PD model family, with its improved path-dependent structure, outperforms the benchmark HAR model family across various loss functions, achieving a notable enhancement in predictive accuracy.
@@
    Our HAR-REQ model passes the MCS test more frequently than the benchmark HAR-REX. Similarly, HAR-PD-REQ and LASSO-HAR-PD-REQ also pass the MCS test more frequently than HAR-REQ. As shown in Tables 13, 14, and 15, the LASSO-HAR-PD-RS model demonstrates the highest predictive accuracy multiple times under the MCS test.
@@ 4.5 Robustness testing
In Section 4.4.1, a rolling window prediction was conducted using the last 600 observations as the prediction sample, following the methodology outlined in Yang et al. (2015). For robustness analysis, a shorter rolling window was selected compared to the out-of-sample forecast, which allowed for a longer prediction interval. To evaluate the model’s performance on actual unknown data, this study employed the first 3,616 observations as the rolling window length and the subsequent 1,000 observations as the prediction window. After data cleaning, the in-sample spanned from January 18, 2005, to November 4, 2019, while the out-of-sample covered the period from November 5, 2019, to December 24, 2023.
Secondly in order to assess the model’s forecasting ability on other data, new data will be used here for forecasting. Section 4.4.1 uses the SSE, while the robustness test uses 970 days of CSI 300 data from January 2, 2020 to December 29, 2023. The data is divided into a 670-day in-sample from January 2, 2020 to October 11, 2022 and a 300-day out-of-sample from October 12, 2022 to December 29, 2023.
Finally, to comprehensively evaluate the superiority of the extended model, additional benchmark models were introduced for comparative analysis beyond the benchmark HAR model family, following the approaches of Qu et al. (2018) and Moreno-Pino and Zohren (2024), who incorporated various traditional benchmark models to demonstrate the performance advantages of their extended models.